# SigmaDock: Untwisting Molecular Docking With Fragment-Based $\operatorname{SE}(3)$ Diffusion

### v1

@@ 3.1 Data and Metrics
Datasets. We use PDBBind(v2020) (Wang et al., 2005), a curated set of 19,443 protein-ligand complexes obtained through crystallography, as our training set. Crucially, we deliberately restrict ourselves to this dataset for fair comparison [^8] It is unfair yet unfortunately common in the literature to compare the same held out set with models trained with larger and more diverse datasets without assessing train-test overlap (Abramson et al., 2024)., isolating any increase in performance obtained in this study to our proposed framework. For validation, we use the well established PoseBusters (Buttenschoen et al., 2024) and Astex (Hartshorn et al., 2007) datasets. PoseBusters(v2) (PB) acts as our temporal-split validation set containing 308
protein-ligand complexes with unseen protein sequences realised from 2021 onwards. The Astex (AX) dataset consists of an additional 85 diverse and highly curated protein-ligand complexes originally designed to faithfully evaluate the quality of protein-ligand docking algorithms.
@@ 3.2 Results
On top of sequence similarity, we stratify the PB set into distinct chemical environments determined by the nature of ligand interactions with additional co-factors (ions, crystallisation aids, natural ligands, or other co-factors). We hypothesise that, since SigmaDock is deliberately designed (for simplicity) to exclude co-factors, higher failure rates should be observed when the true bound pose is realised in conjunction to additional artefacts (co-binding event), as the setup is partially observable. After isolating the protein-ligand pairs for which SigmaDock fails to generate accurate poses [^10] Here we define a failure if the majority of samples generated across $N$ seeds have an RMSD below 2Å., we find this hypothesis to hold true, as per Table 3.2. This result provides additional confidence that SigmaDock does not blindly memorise and hallucinate protein-ligand poses.
Table 1: Ablation results (Top-1 accuracy (%) across the PB set) for different configurations. A,B,C are all re-trained from scratch.
@@
We refer the reader to Appendix I for extended results, and Appendix J for a discussion on the current limitations of our method and future works.
@@ 4 Conclusion
We believe SigmaDock represents a major step forward in the reliability and feasability of deep learning as a promising tool for accelerating drug discovery. Moving away from the torsional parametrisation, we show our proposed $\operatorname{SE}(3)$ diffusion framework is highly competitive with physics-based methods in modelling protein-ligand docking. We extensively lay out the key components of SigmaDock in the Appendices and open-source our codebase [^11] We will release our codebase and model weights after the reviewing process. to proliferate the development of generative models for science. Furthermore, we demonstrate the critical role of principled inductive biases in enabling superior generalisation and data efficiency. From our presented results and analysis, we hope our work encourages rethinking the current trends of simplifying and scaling increasingly larger models (Abramson et al., 2024).
@@ Acknowledgments
AP is funded by EPSRC and AstraZeneca via an iCASE award for a DPhil in Machine Learning. LZ is supported by the EPSRC CDT in Modern Statistics and Statistical Machine Learning (EP/S023151/1). AP thanks Kathryn Giblin for insightful scientific discussions. AP also thanks Yasmin Baba for unconditional support. LZ thanks Jessica Harrison for helpful witticisms in informing (subliminally) our work’s moniker. The authors declare no competing interests.
@@ D.2 Proof of Lemma 1
where gray terms denote cross-edges across fragments that are time-dependent and therefore varying according to the forward-backward diffusion process on SE(3). For clarity, we use the notation $B_{\mathcal{A}}(t)$ to represent the node $B$ of the dihedral $\phi_{\mathcal{A}:BC:\mathcal{D}}$ corresponding to fragment $\mathcal{A}$. This is important because $B$ and $C$ are duplicated (as dummy atoms) in the fragmentation process to preserve torsional bond information. An illustration is shown in Figure 5.
@@ E.1 Data Processing
For a given protein-ligand pair, we define the pocket by including every residue that has at least one atom within a specified cut-off distance $d_{\text{pocket}}$ from any atom in the bound ligand pose. To help de-bias the selection of the pocket we use a stochastic distance $d_{\text{pocket}}=5+\mathcal{N}(0,\sigma_{\text{pocket}})$, where we use a value of $\sigma_{\text{pocket}}=1$ for training and $\sigma_{\text{pocket}}=0.5$ for sampling. If there are multiple ligands present in the SDF file due to the presence of multiple chains, we randomly sample one. We note that we include residues from all protein chains as this is the closest representation of what the ligand sees during its binding event.
@@ I.2 Sample Performance & Efficiency
Table 3: Stratification of the original PDBbind(v2020) vs. AF3 train-test splits on PoseBusters(v2). Sequence similarity split values extracted from AF3 Extended Data Fig. 4c (Abramson et al., 2024). (*) We observe a mismatch between the averaged sequence similarity results (80.1%) vs the reported average performance in Fig. 4e (84%).
@@
$[0,30)$ 109 38 +71+71 72 86 14-14
$[30,95)$ 76 83 7-7 79 80 1-1
$[95,100]$ 123 187 64-64 87 79 +8+8
Total / Avg 308 308 79.9 80.1*
Unexpectedly, AF3’s reported accuracy decreases as test–train similarity increases – contrary to our findings and to prior reports (Škrinjar et al., 2025). Despite AF3 evidently training with substantially higher train–test overlap, our Top-1 accuracy is comparable overall and matches or exceeds AF3 across most of the test set (match on $[30,95)$, surpass on $(95,100]$). For a fair comparison, we deliberately train SigmaDock on a reduced dataset PDBBind(v2020) with $\sim$19k datapoints, and do not model cofactors. Nevertheless, we still achieve AF3-level performance. Contrary to common concerns about the physicochemical coherence of generative deep-learning models for structure prediction, we observe superior generalisation with a compact model (14.9M parameters) and a set of well defined inductive biases. It is therefore reasonable to expect that SigmaDock will further improve with more data, larger models, and added flexibility, and surpass AF3 on more challenging settings such as cross-docking and flexible-receptor protein–ligand docking.

### v2

@@ 3.1 Data and Metrics
Datasets. We use PDBBind(v2020) (Wang et al., 2005), a curated set of 19,443 protein-ligand complexes obtained through crystallography, as our training set. Crucially, we deliberately restrict ourselves to this dataset for fair comparison [^8] It is unfair yet unfortunately common in the literature to compare the same held out set with models trained with larger and more diverse datasets without assessing train-test overlap (Abramson et al., 2024)., isolating any increase in performance obtained in this study to our proposed framework. For validation, we use the well established PoseBusters (Buttenschoen et al., 2024) and Astex (Hartshorn et al., 2007) datasets. PoseBusters(v2) (PB) acts as our temporal-split validation set containing 308 protein-ligand complexes with unseen protein sequences realised from 2021 onwards. The Astex (AX) dataset consists of an additional 85 diverse and highly curated protein-ligand complexes originally designed to faithfully evaluate the quality of protein-ligand docking algorithms.
@@ 3.2 Results
On top of sequence similarity, we stratify the PB set into distinct chemical environments determined by the nature of ligand interactions with additional co-factors (ions, crystallisation aids, natural ligands, or other co-factors). We hypothesise that, since SigmaDock is deliberately designed (for simplicity) to exclude co-factors, higher failure rates should be observed when the true bound pose is realised in conjunction to additional artefacts (co-binding event), as the setup is partially observable. After isolating the protein-ligand pairs for which SigmaDock fails to generate accurate poses [^10] Here we define a failure if the majority of samples generated across $N$ seeds have an RMSD above 2Å., we find this hypothesis to hold true, as per Table 3.2. This result provides additional confidence that SigmaDock does not blindly memorise and hallucinate protein-ligand poses.
Table 1: Ablation results (Top-1 accuracy (%) across the PB set) for different configurations. A-C are re-trained from scratch;   (*): default.
@@
[…]
Table 3:   Sensitivity to pocket definition (PB set). Pocket diameter $\varnothing_{\text{pocket}}$ is the maximum pairwise distance between $C_{\alpha}$’s across the selected $N_{\textbf{res}}$ residues; we report dataset means.
Metric $||$ $d_{0}$(Å) 4 5 6 7
$\varnothing_{\text{pocket}}$ (Å) 20.9 22.4 24.2 26.2
$N_{\text{res}}$ 15.6 20.3 26.5 37.8
RMSD << 2 80.5 81.5 78.3 69.8
PB-Val. 80.2 80.5 77.3 68.2
Table 4:   Per-sequence-similarity comparison between SigmaDock (left) and AF3 (right) on the PB set. Values for AF3 are extracted from Extended Data 4c (Abramson et al., 2024).
Seq. Sim. (%) Count PB-Val.
$[0,30)$ 109 $|$ 38 72 $|$ 87
$[30,95)$ 76 $|$ 83 79 $|$ 82
$[95,100]$ 123 $|$ 187 87 $|$ 78
Total / Avg. 308 $|$ 308 79.9 $|$ 80.2
We refer the reader to Appendix I for extended results, and Appendix J for a   detailed discussion on the current limitations of our method and future work.
@@ 4 Conclusion
We believe SigmaDock represents a major step forward in the reliability and feasibility of deep learning as a promising tool for accelerating drug discovery. Moving away from torsional parametrisation, our proposed $\operatorname{SE}(3)^{m}$   fragment-space Riemannian diffusion model is, to our knowledge, the first generative method trained on the intended PB split to surpass classical dockers on the re-docking task. We extensively lay out the key components of our framework in the Appendices and open-source our codebase to proliferate   reproducibility and further development, and we view extensions to flexible-receptor docking and co-folding as natural next steps to make SigmaDock a more general and practical tool. We demonstrate the critical role of principled inductive biases in enabling superior generalisation and data efficiency, and hope our work encourages rethinking progress   on geometry and conditioning, as opposed to relying on scale alone (Abramson et al., 2024).
@@ Acknowledgments
AP is funded by EPSRC and AstraZeneca via an iCASE award for a DPhil in Machine Learning. LZ is supported by the EPSRC CDT in Modern Statistics and Statistical Machine Learning (EP/S023151/1). AP thanks Kathryn Giblin for insightful scientific discussions, as well as Jochem Nelen and other PhD students in the OPIG and OxCSML groups for valuable exchanges. AP also thanks Yasmin Baba for her unwavering support and for compelling him to take a rare break from research, during which the core idea of this work emerged. LZ thanks Jessica Harrison for helpful witticisms in informing (subliminally) our work’s moniker. The authors declare no competing interests.
Reproducibility Statement & Code Availability
All training and evaluation data used in this study are publicly available. The full codebase for training, sampling, and evaluation used to generate the reported results will be released open source upon publication at github.com/alvaroprat97/sigmadock.
@@ D.2 Proof of Lemma 1
where gray terms denote cross-edges across fragments that are time-dependent and therefore varying according to the forward-backward diffusion process on $\operatorname{SE}(3)$. For clarity, we use the notation $B_{\mathcal{A}}(t)$ to represent the node $B$ of the dihedral $\phi_{\mathcal{A}:BC:\mathcal{D}}$ corresponding to fragment $\mathcal{A}$. This is important because $B$ and $C$ are duplicated (as dummy atoms) in the fragmentation process to preserve torsional bond information. An illustration is shown in Figure 5.
@@ D.4 FR3D
Degrees of freedom (DoF)
Let the ligand be partitioned into $m$ rigid fragments; the global pose is $z\in\mathrm{SE}(3)^{m}$ with $6m$ DoFs (3 rotation, 3 translation per fragment), which is far smaller than the $3N$ DoFs (coordinates) of all-atom point clouds ($N\gg m$). Naïve fragmentation by cutting all $k$ rotatable bonds yields $\hat{m}=k+1$ fragments, i.e., $6\hat{m}$ DoFs, which exceeds the torsional parameterisation ($k+6$). Our FR3D merging reduces $m$ empirically to $\approx\tfrac{2}{3}(k{+}1)$, lowering the raw coordinate count. Across each cut bond we introduce two $\operatorname{SE}(3)$-invariant triangulation distances (Sec. 2.2.3; Lemma 1) that fix adjacent bond angles and the torsional bond length while leaving the dihedral free. Denote by $e$ the number of cut bonds (edges) between fragments and stack all distance constraints into $c(z)\in\mathbb{R}^{2e}$. The hard-constrained manifold
$\mathcal{M}_{\mathrm{frag}}=\{\,z\in\mathrm{SE}(3)^{m}:\ c(z)=0\,\}$
has tangent dimension
$\dim\mathcal{M}_{\mathrm{frag}}\;=\;6m\;-\;\operatorname{rank}\!\big(J_{c}(z)\big).$
For a tree fragment graph (e=m1e=m{-}1), two non-degenerate point-to-point distance constraints across a joint generically remove $5$ relative DoF (leaving a single revolute DoF about the bond axis), so $\operatorname{rank}(J_{c})=5(m{-}1)$ and
$\dim\mathcal{M}_{\mathrm{frag}}\;=\;6m-5(m{-}1)\;=\;m+5.$
Under naïve fragmentation (m=k+1m=k{+}1) the hard (torsional) lower bound is $k+6$. With FR3D, each merge removes one cut and we drop one triangulation distance at the adjacent bond to avoid over-constraining the dihedral across adjacent fragments. Hence, the effective lower bound remains $k+6$ after merging, while the upper bound tightens to $6m$ such that $k+6\leq\text{DoF}_{\text{eff}}\leq 6m$.
Soft vs. hard constraints. In practice, triangulation is injected as soft, $\operatorname{SE}(3)$-invariant edge features $c_{t}$ at each denoising step; no hard equalities are imposed. Thus trajectories concentrate near $\mathcal{M}_{\mathrm{frag}}$ as $t\to 0$, and the effective DoF lies between $k+6$ and $6m$. Consistent with Lemma 1, the only local DoF around a cut bond as $t\to 0$ is the dihedral $\Delta\phi_{ABCD}$. Empirically, removing these conditioning features reduces PB-validity by $12.8\%$ (Table 3.2, Config A).
@@ E.1 Data Processing
For a given protein-ligand pair, we define the pocket by including every residue that has at least one atom within a specified cut-off distance $d_{0}$ from any atom in the bound ligand pose. To help de-bias the selection of the pocket we use a stochastic cutoff $d_{r}=d_{0}+\mathcal{N}(0,\sigma_{r})$, where we use default values of $\sigma_{r}=1,d_{0}=5$ for training and $\sigma_{r}=0.5$ for sampling (all in Å). If there are multiple ligands present in the SDF file due to the presence of multiple chains, we randomly sample one. We note that we include residues from all protein chains as this is the closest representation of what the ligand sees during its binding event.
@@ I.2 Sample Performance & Efficiency
Table 5: Stratification of the original PDBbind(v2020) vs. AF3 train-test splits on PoseBusters(v2). Sequence similarity split values extracted from AF3 Extended Data Fig. 4c (Abramson et al., 2024). (*) We observe a mismatch between the averaged sequence similarity results (80.2%) vs the reported average performance in AF3’s Extended Data Fig. 4e (84.4%).
@@
$[0,30)$ 109 38 +71+71 72 87 16-16
$[30,95)$ 76 83 7-7 79 82 3-3
$[95,100]$ 123 187 64-64 87 78 +9+9
Total / Avg 308 308 79.9 80.2*
Unexpectedly, AF3’s reported accuracy decreases as test–train similarity increases – contrary to our findings and to prior reports (Škrinjar et al., 2025). Despite AF3 evidently training with substantially higher train–test overlap, our Top-1 accuracy is comparable overall and matches or exceeds AF3 across most of the test set (match on $[30,95)$, surpass on $(95,100]$). For a fair comparison, we deliberately train SigmaDock on a reduced dataset PDBBind(v2020) with $\sim$19k datapoints, and do not model cofactors. Nevertheless, we still achieve AF3-level performance. Contrary to common concerns about the physicochemical coherence of generative deep-learning models for structure prediction, we observe superior generalisation with a compact model (14.9M vs. $\sim$500M parameters for AF3) and a set of well defined inductive biases. It is therefore reasonable to expect that SigmaDock will further improve with more data, larger models, and added flexibility, and surpass AF3 on more challenging settings such as cross-docking and flexible-receptor protein–ligand docking.